# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ 1.2 AI techniques
This study investigates AI-based techniques for processing unstructured text data to analyse fraud. Much of this text data, such as news articles, research papers, government reports, books, social media posts (such as tweets and Facebook comments), communications (such as emails, SMS messages, and chat logs), and web content (such as reviews on online marketplaces, travel and hospitality platforms, and comments on video sharing platforms), is inherently unstructured. Statista reported the global data created, captured, copied, and consumed is 64.2 zettabytes in 2020, and it is expected to exceed 180 zettabytes by 2025 Statista . With each new digital platform or communication channel, this data is increasing. Most of this created data is unstructured text data that provides opportunities for understanding human behaviour, habits, opinions and experiences. It contains information about users’ experiences, events, themes, opinions and sentiments that can be important for deriving meaningful insight from their experience related to fraudulent activities. Manual traditional data analysis techniques like keyword searches and coding of themes, are often limited and unachievable to extract meaningful insights, making advanced computer-driven automated techniques necessary.
However, this data often demonstrates significant challenges due to the diversity of natural (human) language. This includes dealing with noise, which includes irrelevant data, a wide array of linguistic variations of human language due to regional or cultural nuances, use of slang or jargon, abbreviations, spelling errors, typos and grammatical mistakes, which often pose challenges for analysing text data efficiently. NLP techniques were designed to effectively understand the structure (syntax) and comprehend (semantic) spoken and written human language the way humans do. Advancements in AI, including machine and deep learning, along with improvements in technology (such as increased computing power) and software (such as the availability of programming tools and libraries), have significantly improved the ability to process and understand large volumes of unstructured text. These tools have been widely used in many fields from sales and marketing to spam detection. The process of collecting, pre-processing, and training AI-based models using text data often involves the pipeline shown in Figure1:
@@
  • Problem statement: Using domain knowledge, a suitable research question is formulated for AI to address. This could be a classification problem (e.g. to classify text into a number of categories), or explanatory analysis involving identifying patterns within a text.
  • Data preparation: AI-based models require the collection of appropriate data to train models to analyse the research question. Often, the acquired data comes as unstructured data and requires cleaning and pre-processing. The data collection involves the collection of appropriate data and building a corpus. A corpus is a collection of large and structured sets of documents (e.g. emails, news articles, social media posts or transcripts). The data cleaning and pre-processing involves removing unwanted or redundant data to reduce the noise in the data. This may include removing duplicates or incomplete entries, removing symbols, punctuation, numbers, stop-words, converting acronyms to full words, and handling non-English words and slang or a particular jargon. Further pre-processing may involve text normalisation techniques like stemming or lemmatisation to reduce words to their root or base form to improve the accuracy of text analysis.
  • Feature engineering and selection: Feature engineering involves preparing data for machine learning models, and it involves the extraction and selection of predictive features in supervised learning) or find patterns and structures in unlabeled data in unsupervised learning. This task requires using domain knowledge to develop and select appropriate features. Common text features often used are n-grams (sequences of n consecutive words); Term Frequency-Inverse Document Frequency (TF-IDF) (a statistical method that weighs the importance of a word/term to a document within a corpus) matrix; sentiments and emotions present; lexical features (e.g. presence of certain words, Keyword-in-Context and lexical diversity); syntactic features (e.g. part-of-speech tags); semantic features (e.g. entities mentioned, word-embedding); readability scores; structural features (e.g. length of the text, number of paragraphs); and domain-specific features (e.g. presence of certain specialised terms).
  • AI technique/algorithm selection: This step involves selecting an appropriate AI algorithm for building the AI-based model. The text tasks often involve two main categories of AI-based models: supervised and unsupervised machine learning. The choice of the algorithm will depend on the learning, and type of AI required. Supervised machine learning algorithms are often used for text classification problems. The learning algorithm is provided with input features (training data) and labels (discrete outputs). The goal of the supervised machine learning algorithm is to map input features to discreet outputs. Traditional supervised machine learning algorithms include Logistic Regression, Naive Bayes, decision trees, Random Forest (RF) (multiple decision trees), Support Vector Machines (SVM) and K-Nearest Neighbors (KNN). New techniques of supervised machine learning include neural networks and deep learning-based models. Unsupervised machine learning models, often used in exploratory data analysis, involve working with unlabelled data to discover hidden patterns and themes. Unsupervised machine learning algorithms including clustering techniques using algorithms like K-means, hierarchical clustering and DBSCAN (Density-Based Spatial Clustering); and topic modelling using algorithms like Latent Dirichlet Allocation and Latent Semantic Analysis (LSA).
@@
  • Model training: Often machine learning algorithms will have parameters that need to be tuned before learning begins, these are known as hyperparameters.The tuning process involves re-training the model multiple times using different values for these hyperparameters and selecting the best combination of values based on model performance on a metric of interest. In the case of supervised machine learning, the hyperparameters might be tuned using model performance on different “folds” of the data in an approach known as cross-validation, with some randomly-selected proportion of the data kept separate from the training data, known as a test dataset, for final model evaluation, giving the best available indication of how the model is likely to perform on new, unseen data. In the case of unsupervised modelling, heuristics will be used to identify the optimal number of clusters or topic modelling.
@@
  • Model evaluation: The performance of the model needs to be evaluated. In the case of supervised modelling, this will involve measuring the performance of the model on the test data. The classic supervised machine learning algorithms can be evaluated using performance metrics such as confusion matrix (Figure 2), accuracy, precision, recall, F1-score, sensitivity, specificity, Receiver Operating Characteristic (ROC) curve, and the Area Under the Curve (AUC) curve:
@@
    True Positives (TP): The model correctly predicts positive class (e.g., those that were classified as scam.)
    True Negatives (TN): The model correctly predicts negative class (e.g., those classified as not-scam).
    False Positives (FP): The model incorrectly predicts the positive class but it is not (e.g. not-scam is predicted as a scam).
    False Negatives (FN): The model incorrectly predicts the negative class but it is not (e.g. scam predicted as not scam).
    Sensitivity and specificity are two other performance measures. Sensitivity is the same as recall, or true positive rate that captures the model’s ability to correctly identify positive class (i.e. scam cases):
@@
    Specificity, also known as false positive rate (FPR) measures the proportion of true negatives, and it captures the model’s ability to correctly identify negative class (i.e. not-scam cases):
@@
    The ROC curve is a graph illustrating the performance of one or more binary classifiers. It plots the sensitivity against the 1 - 1-specificity for various thresholds. The AUC is calculated as the area under the ROC curve.
  • Deploy model: Once the models are working well they can be deployed. When considering deployment of the model, one must address questions regarding why others should trust the model, how the model arrived at its conclusions and usability, and carefully assess the ethical implications of AI to ensure its suitability for deployment and that it is not biased.
In cases of unsupervised machine learning models, due to a lack of ground truth labels, the performance of the model evaluation may involve subjective interpretation to interpret the outputs (e.g. clusters or topics) generated by the model.
@@ 1.3 Advanced NLP techniques
This section provides a brief introduction to advanced NLP techniques, aiming to familiarise the reader with these concepts as they are later referred to during the SLR findings.
@@
Word embeddings is a technique in NLP that encodes words as vectors of real numbers to capture their similarities. Words that are closer together in the vector space are expected to have similar meanings or relationships. Two of the widely used word embedding techniques are Word2Vec and GloVe. Word2Vec uses a simple neural network trained on large text datasets iteratively to predict either context words or target words. Word2Vec uses two approaches DBLP:journals/corr/Rong14 : Continuous Bag of Words (CBOW) predicts the target word based on its surrounding context words, whereas Skip-gram predicts surrounding context words based on a given target word. In the sentence ’The quick brown fox jumps over the lazy dog’, if ’fox’ is used as the target word, the CBOW model uses ’The’, ’quick’, ’brown’, ’jumps’, ’over’, ’the’, ’lazy’, and ’dog’ as context and predicts the word ’fox’. In Skip-gram, ’fox’ is used to predict the surrounding words like ’The’, ’quick’, ’brown’, ’jumps’, ’over’, ’the’, ’lazy’, and ’dog’. Global Vectors for Word Representation (GloVe glove learns the vector representation of words using global word-word co-occurrence statistics obtained from the training data to show the semantic relationships between words.
Word embeddings are often used by Large Language Models to comprehend and respond to language.
@@
Large Language Models(LLMs) zhao2023surveylargelanguagemodels are advanced NLP tools trained on billions of words from a wide variety of sources such as newspapers, books and websites, which are designed to perform complex tasks such as translations, summarisation and the performance of human-like conversational abilities. Most LLMs are developed using a transformer-based architecture (transformers) DBLP:journals/corr/VaswaniSPUJGKP17 and they have billions of parameters used for training. Transformers are a type of deep-learning neural network model, and they are more efficient compared with predecessor state-of-the-art models which were based on Recurrent Neural Network (RNN). Transformers use a complicated architecture that has encoder and decoder layers to understand sequences of words and provide an output DBLP:journals/corr/VaswaniSPUJGKP17 . While the encoder layer processes input text data, extracting hierarchical representations through mechanisms like self-attention, the decoder layers generate output sequences based on the input received from the encoder. Transformer-based LLMs include GPT models including GPT3 (Generative Pre-trained Transformer 3), GPT-4 and GPT-4o by OpenAI. Both GPT-4 and GPT-4o are multimodal models that accept text and image and produce text chatgpt2024 . Other transformers include Bidirectional Encoder Representations from Transformers (BERT), and its smaller and lighter version of DistilBERT which is designed for applications where computational resources might be limited. BERT and DistilBERT are also designed to understand context in language processing and are suitable for NLP tasks like text classification, answering questions and named entity recognition. The difference between models like BERT and GPT is the way their architecture is designed and the objectives of learning. BERT uses only the encoder of the transformer architecture and processes text bidirectionally by considering both preceding and following words. On the other hand, GPT models use only a decoder of the transformer architecture and process text in an unidirectional manner by processing text based only on preceding words.
@@
LLMs can assist in analysing large amounts of text data and identify patterns automatically which can be useful when dealing with fraud and other crime-related data, they can be misused by criminals to generate content for fraudulent activities. Generative AI refers to AI techniques that can be used to generate new text, audio, images and video that closely resembles human-generated content. LLMs have been successfully applied in various areas of human communications including chatbots in customer support systems by generating human-like text, content generation and performing language translation. However, this same technology can be exploited by criminals to create content for deception including fake websites, targeted phishing emails and scam advertisements to deceive potential victims.
1.4 Research approach
@@
Although there are some literature reviews related to the application of AI for fraud and crime, to our best knowledge there are currently no SLRs that attempt to understand the state-of-the-art towards detecting online fraud in general. Overall, the literature reviews that we found discuss the state-of-the art towards detecting specific online fraud or scams, e.g., credit card fraud alone. In this work, we aim to understand whether there are universal AI methodologies that attempt to detect online fraud, focusing on textual data.
Search String
("Online Fraud*" OR "scam*") AND (("machine learning")
OR "NLP" OR ("natural language processing") OR "classifier"
OR ("Large Language Models") OR "LLM"
OR ("Generative Artificial Intelligence") OR "GenAI" OR "GAI")
@@
Table 1: Query for the literature selection in various academic libraries.
2 Systematic Review
@@
Systematic reviews differ from traditional literature reviews as they aim to identify all relevant studies that address a set of research questions using a methodology that is structured and can be replicated nightingale2009guide .
@@ 2.1 Methods
The following methodology was employed to conduct the SLR, and address the selection process for identifying relevant publications and to avoid biases.
@@ 2.1.1 Protocol
For this SLR we follow the Preferred Reporting Items for Systematic Reviews and Meta-Analysis extensions for Scoping Reviews (PRISMA-ScR), as proposed by Moher et al. moher2010preferred .
2.1.2 Eligibility
This review focuses on studies that use AI-based models, specifically NLP models, including Machine Learning (ML) and Deep Learning (DL) techniques using text data. For example, studies that employ AI to detect fraudulent bank accounts, fraudulent credit card transactions, or fraudulent networks of users online, are not to be included.
To be considered for inclusion in the SLR, we used the following eligibility criteria:
  • Peer-Reviewed Studies: We focused on peer-reviewed studies published in English between 2019 and March 2024. Our search was restricted to academic records found in journals and conference proceedings. We excluded theses, legal documents, patents, and citations from these studies.
  • Grey Literature:To capture the latest AI-based models, we also included grey literature, specifically pre-prints from ArXiv published in 2023 that have not yet been incorporated into conference proceedings or book chapters.
  • Search Strategy: Table 1 shows the search strings used to detect related literature in ACM Library, ProQuest, Web of Science, IEEE Xplore, ArXiv, and Google Scholar. The authors finalized this string after trying various searches in these libraries.
  • Scope and Focus: Studies must address fraud performed online and use AI-based methodologies for analysing online fraudulent activities. The focus is on studies using textual data, whether from scammers, victims, or victim reports, to understand, detect, or analyse online fraud activities. Our goal is to understand what is the state-of-the-art on models that are designed to prevent and detecting scams (tricking the victim) before the victim gets defrauded which will lead them to monetary losses. Studies that analyse money transactions, credit card transactions, and cryptocurrency transactions are excluded, from this review, as they do not use text data.
  • AI-Methodology: Papers must include a methodology section or a similar section where the authors discuss their AI implementation and fine-tuning along with the accuracy of their model. Finally, we consider studies published after 2019.
  • Publication timeframe: Papers published between 2019 and March 2024 are included in this review. This period was selected to manage the overwhelmingly large volume of papers on online fraud and to align with our available resources. Also, we believe that studies conducted before 2019 are less likely to reflect recent advancements in AI methods. Given the rapid evolution of AI-based models, focusing on the 2019-March 2024 period ensures the inclusion of the most up-to-date and relevant research.
@@ 2.1.2 Eligibility
Title The title of the manuscript Detecting Phishing URLs Using Deep Learning
@@
Data Quantity The amount of data used for the analysis 5000 phishing URLs and 5000 legitimate URLs
Models Used All models the authors experimented with Random Forest, Linear Regression, LDA, W2V
@@
Model Stats All performance metrics A=0.95, P=0.80, R=0.9, AUC=0.89

### v2

@@ 2.2 AI techniques
This study investigates AI-based techniques for processing unstructured text data to analyse fraud. Much of this text data, like news articles, research papers, government reports, books, social media posts (tweets and Facebook comments), communications (such as emails, SMS messages, and chat logs), and web content (reviews on online marketplaces, travel and hospitality platforms, and comments on video sharing platforms), is inherently unstructured. Statista Statista estimated that the global open data that is accessible on the entire Web was 64.2 zettabytes in 2020, and it is expected to exceed 180 zettabytes by 2025. With each new digital platform or communication channel, this data is increasing. Most of the data created is unstructured text that provides opportunities for understanding human behaviour, habits, opinions and experiences. It contains information about users’ experiences, events, themes, opinions, and sentiments that can be important for deriving meaningful insight from their experience related to fraudulent activities. Manual traditional data analysis techniques, like keyword searches and the coding of themes, are often limited, and the extraction of meaningful insights at scale is unachievable, making advanced computer-driven automated techniques necessary.
However, there are often significant challenges associated with the analysis of this data due to the diversity of natural (human) language used. This includes dealing with noise (irrelevant or useless data), a wide array of linguistic variations of human language due to regional or cultural nuances, the use of slang or jargon, abbreviations, spelling errors, typos and grammatical mistakes, which often pose challenges for the efficient analysis of text data. NLP techniques were designed to effectively understand the structure (syntax) and comprehend (semantic) spoken and written human language the way humans do. Advancements in AI, including machine and deep learning, along with improvements in technology (such as increased computing power) and software (such as the availability of programming tools and libraries), have significantly improved the ability to process and understand large volumes of unstructured text. These tools have been widely used in many fields, from sales and marketing to spam detection. The process of collecting, pre-processing, and training AI-based models using text data often involves the pipeline shown in Figure 1:
@@
  • Problem statement: Using domain knowledge, a suitable research question is formulated for AI to address. This could be a classification problem (e.g. to classify text into a number of categories), or explanatory analysis involving the identification of patterns within a text.
  • Data preparation: AI-based models require the collection of appropriate data towards the building of a corpus (a collection of structured sets of documents such as emails, news articles, social media posts, or transcripts) used to train models to analyse the research question. Often, the acquired data comes as unstructured data and requires cleaning and pre-processing. The data cleaning and pre-processing involve removing unwanted or redundant data to reduce the noise in the data. This may include removing duplicates or incomplete entries, symbols, punctuations, numbers, stop-words, converting acronyms to full words, and handling non-English words, slang, or jargon. Further pre-processing may involve text normalisation techniques like stemming or lemmatisation to reduce words to their root or base form to improve the accuracy of text analysis.
  • Feature engineering and selection: Feature engineering involves preparing data for machine learning models. It consists of extracting and selecting predictive features in supervised learning or finding patterns in unlabeled data in unsupervised learning. This task requires using domain knowledge to develop and select appropriate features. Common text features often used are n-grams (sequences of n consecutive words); Term Frequency-Inverse Document Frequency (TF-IDF) (a statistical method that weights the importance of a word/term in a document within a corpus) matrix; sentiments and emotions present; lexical features (e.g. presence of certain words, Keyword-in-Context and lexical diversity); syntactic features (e.g. Part-of-Speech tags); semantic features (e.g. entities mentioned, word-embedding); readability scores; structural features (e.g. length of the text, number of paragraphs); and domain-specific features (e.g. presence of specialised terms).
  • AI technique/algorithm selection: This step involves selecting an appropriate AI algorithm for building the AI-based model. Tasks associated with text often involve two main categories of AI-based models: supervised and unsupervised machine learning. The choice of the algorithm will depend on the learning, and type of AI required. Supervised machine learning algorithms are often used for text classification problems. The learning algorithm is fed with input features (training data) and labels (discrete outputs). The supervised machine learning algorithm aims to map input features to discreet outputs. Traditional supervised machine learning algorithms include Logistic Regression (LR), Naive Bayes (NB), Decision Trees (DT), Random Forest (RF - multiple DTs), Support Vector Machines (SVM), and K-Nearest Neighbors (KNN). New supervised machine learning techniques include neural networks (NN) and deep learning-based models. Unsupervised machine learning models, often used in exploratory data analysis, involve working with unlabelled data to discover hidden patterns and themes. Unsupervised machine learning algorithms include clustering techniques using algorithms like K-means, hierarchical clustering and Density-Based Spatial Clustering (DBSCAN); and topic modelling, achieved by algorithms like Latent Dirichlet Allocation (LDA) and Latent Semantic Analysis (LSA).
@@
  • Model training: Often, machine learning algorithms will have parameters that need to be tuned before learning begins, known as hyperparameters. The tuning process involves re-training the model multiple times using different values for these hyperparameters and selecting the best combination of values based on model performance on a metric of interest. In the case of supervised machine learning, the hyperparameters might be tuned using model performance on different “folds” of the data in an approach known as cross-validation. With this approach, a randomly selected proportion of the data is kept separate from the training data, and used for final model evaluation. This approach provides the best indication of how the model will likely perform on new, unseen data. In the case of unsupervised modelling, heuristics are used to identify the optimal number of clusters or topic modelling.
@@
  • Model evaluation: The model’s performance needs to be evaluated. In the case of supervised modelling, this will involve measuring the model’s performance on the test data. The classic supervised machine learning algorithms can be evaluated using performance metrics such as a confusion matrix (Figure 2), accuracy, precision, recall, F1-score, sensitivity, specificity, Receiver Operating Characteristic (ROC) curve, and the Area Under the Curve (AUC) curve:
@@
    True Positives (TP): The model correctly predicts a positive class (e.g., those that were classified as scam.)
    True Negatives (TN): The model correctly predicts a negative class (e.g., those classified as not-scam).
    False Positives (FP): The model incorrectly predicts the positive class (e.g. not-scam is predicted as a scam).
    False Negatives (FN): The model incorrectly predicts the negative class (e.g. scam predicted as not scam).
    Sensitivity is the same as recall or the true positive rate, and it captures the model’s ability to identify the positive class (i.e. scam cases) correctly:
@@
    Specificity, also known as false positive rate (FPR), measures the proportion of true negatives, and it captures the model’s ability to identify negative class (i.e. not-scam cases) correctly:
@@
    The ROC curve illustrates the performance of one or more binary classifiers. It plots the sensitivity against the 1-specificity for various thresholds. The AUC is calculated as the area under the ROC curve.
  • Deploy model: Once the models work well, they can be deployed. When considering deployment of the model, one must address questions regarding why others should trust the model, how the model arrived at its conclusions and usability, and carefully assess the ethical implications of AI to ensure its suitability for deployment and that it is not biased.
In unsupervised machine learning models, due to a lack of ground truth labels, the performance of the model evaluation may involve subjective interpretation to interpret the outputs (e.g. clusters or topics) generated by the model.
@@ 2.3 Advanced NLP techniques
This section briefly introduces advanced NLP techniques, aiming to familiarise the reader with these concepts as they are later referred to during the SLR findings.
Word embeddings is an important technique in NLP that involves encoding words as vectors of real numbers that are designed to capture their similarities.
@@
Words closer together in the vector space are expected to have similar meanings or relationships. Two of the widely used word embedding techniques are Word2Vec and GloVe. Word2Vec uses a simple neural network trained on large text datasets iteratively to predict either context words or target words. Word2Vec uses two approaches DBLP:journals/corr/Rong14 : Continuous Bag of Words (CBOW) predicts the target word based on its surrounding context words, whereas Skip-gram predicts surrounding context words based on a given target word. In the sentence ‘The quick brown fox jumps over the lazy dog’, if ‘fox’ is used as the target word, the CBOW model uses ‘The’, ‘quick’, ‘brown’, ‘jumps’, ‘over’, ‘the’, ‘lazy’, and ‘dog’ as context and predicts the word ‘fox.’ In Skip-gram, ‘fox’ is used to predict the surrounding words like ‘The’, ‘quick’, ‘brown’, ‘jumps’, ‘over’, ‘the’, ‘lazy’, and ‘dog’. Global Vectors for Word Representation (GloVe) glove learns the vector representation of words using global word-word co-occurrence statistics obtained from the training data to show the semantic relationships between words.
Large Language Models use word embeddings to generate responses to natural language inputs.
@@
Large Language Models (LLMs) zhao2023surveylargelanguagemodels are advanced NLP tools trained on billions of words from a wide variety of sources and are designed to perform complex tasks like translations, summarisation and the performance of human-like conversational abilities. Most LLMs are developed using a transformer-based architecture (transformers) DBLP:journals/corr/VaswaniSPUJGKP17 , and billions of parameters are used for training. Transformers are a type of deep-learning neural network model, and they are more efficient compared with predecessor state-of-the-art models based on Recurrent Neural Networks (RNN). Transformers use a complicated architecture with encoder and decoder layers to understand sequences of words and provide an output DBLP:journals/corr/VaswaniSPUJGKP17 . While the encoder layer processes input text data, extracting hierarchical representations through mechanisms like self-attention, the decoder layers generate output sequences based on the input received from the encoder. Transformer-based LLMs include GPT models like Generative Pre-trained Transformer 3 (GPT-3), GPT-4 and GPT-4o developed by OpenAI. GPT-4 and GPT-4o are multimodal models that accept text and image and produce text chatgpt2024 . Other transformers include Bidirectional Encoder Representations from Transformers (BERT) and its smaller and lighter version of DistilBERT, designed for applications with limited computational resources. BERT and DistilBERT are also designed to understand context in language processing and are suitable for NLP tasks like text classification, answering questions, and named entity recognition. The difference between models like BERT and GPT is the way their architecture is designed and the intended learning objectives.
@@
LLMs can assist in analysing large amounts of text data and identify patterns automatically, which can be helpful when dealing with fraud and other crime-related data. LLMs have been successfully applied in various areas of human communications, including chatbots in customer support systems, by generating human-like text, content generation, and performing language translation. Generative AI (GenNAI or GAI) refers to AI techniques that create new text, audio, images and video that closely resemble human-generated content. On the other hand, criminals can misuse these resources to generate content for fraudulent activities, such as fake websites, targeted phishing emails, and scam advertisements, to deceive potential victims.
2.4 The Use of AI in Fraud Detection
@@
Although some literature reviews explore the application of AI for fraud and crime, to the best of our knowledge, no reviews currently aim to understand the state-of-the-art in detecting online fraud in general. The literature reviews we found, discuss the detection of specific online fraud or scams, such as credit card fraud cherif2023credit and phishing SMS barrera2023literature , among others.
In more detail, our preliminary analysis of literature reviews finds that specific AI models work best towards detecting specific types of fraud (e.g., phishing URLs, smishing, etc.), as researchers perform literature reviews to analyse specific offences and not analyse the general task of fraud overall. In addition, a single/universal model does not perform well at classifying various types of fraud. Hence, researchers must constantly develop and update their trained models to detect specific fraud types. In this work, we aim to understand whether there are universal AI methodologies that attempt to detect online fraud, in general, focusing on textual data.
Search String Library Notes
("Online Fraud*" OR "scam*") ACM All text
AND (("machine learning") ProQuest All text, Journals, Conferences
OR "NLP" OR ("natural language processing") Web of Science Topic
OR "classifier" OR ("Large Language Models") OR "LLM" IEEE Xplore Abstract, Journals, Conferences
OR ("Generative Artificial Intelligence") arXiv Abstract, Computer Science, Jan 2023-Mar 2024
OR "GenAI" OR "GAI") Google Scholar All text, Review articles, 2023-2024
@@
Table 1: Search query for the literature selection in various academic libraries. Notes depict the advanced search filters applied to each library.
Research Questions:
  • RQ2: What are the data sources researchers use to analyse online fraud?
  • RQ3: How do researchers evaluate their AI models?
  • RQ4: What are the most popular fraud activities that researchers studied?
Although a wide number of studies have explored the application of AI for fraud detection and other types of cybercrime, we are unaware of any systematic literature reviews that have examined the application of AI models using text data. This SLR focuses on AI-based models that study textual data to detect and gain insights about online fraud. Thus, this study identifies NLP models used to detect online fraud.
3 Systematic Review Methodology
@@
Systematic reviews differ from traditional literature reviews as they aim to identify all relevant studies that address a set of research questions using a structured methodology that can be replicated nightingale2009guide .
@@ 3.1 Methods
We use the following methodology to conduct the SLR and address the selection process to identify relevant publications and avoid biases.
@@ 3.1.1 Protocol
We followed the Preferred Reporting Items for Systematic Reviews and Meta-Analysis extensions for Scoping Reviews (PRISMA), as proposed by Moher et al. moher2010preferred . In a nutshell, this provides a comprehensive framework for conducting and reporting systematic reviews and meta-analyses. The process includes a checklist and flow diagram to ensure transparency, reproducibility, and rigour in summarizing research evidence, to improve the quality of reviews in various fields and to standardise how a literature review should be reported.
@@
Title The title of the manuscript Detecting Phishing URLs Using NLP
@@
Data Quantity The amount of data used for the analysis 100 phishing URLs, 100 legitimate URLs
Models Used All models the authors experimented with RF, LDA, W2V
@@
Model Stats All performance metrics A=0.95, P=0.81, R=0.9, AUC=0.89
@@
3.1.2 Eligibility
This review focuses on studies that use AI-based models, specifically NLP models, including Machine Learning (ML) and Deep Learning (DL) techniques using text data. For example, studies that employ AI to detect fraudulent bank accounts, fraudulent credit card transactions, or fraudulent networks of users online were out of scope. For a study to be considered for inclusion in the SLR, we used the following eligibility criteria:
  • Peer-Reviewed Studies: We focused on peer-reviewed studies published in English between January 2019 and March 2024. Our search was restricted to academic records found in journals and conference proceedings. We excluded theses, legal documents, patents, and citations.
  • Grey Literature: To capture the latest AI-based models, we also included grey literature, specifically pre-prints from ArXiv published between January 2023 and March 2024 that have not yet been incorporated into conference proceedings or academic journals. We also conducted searches on Google Scholar between January 2023 and March 2024.
  • Search Strategy: Table 1 shows the search string used to query related literature in ACM Library, ProQuest, Web of Science, IEEE Xplore, arXiv, and Google Scholar. This was finalised after trying various searches in these libraries. Due to the different functionalities of each library, we had to adjust our search accordingly: for ACM library, we queried our search string across the entire text and adjusted the time range; for ProQuest, we queried our search string across the entire text, adjusted the time range, included only papers from conferences and journals, included only full text and peer-reviewed results, and filtered the subjects to exclude medical terms; for Web of Science we queried our search string on the topic (title, abstract, and keywords) of the paper and adjusted the time range; for IEEE Xplore, we queried our search string on the abstract of the papers, adjusted the time range, and filtered results for papers published in conference proceedings and journals; for arXiv, we queried our search string on the abstract of papers, published in computer science between January 2023 and March 2024; and finally, for Google Scholar, we queried our search string and filtered our results on review articles published between 2023 and 2024. The adjustments mentioned above were implemented to better capture literature related to the scope of our study, and a consensus was reached after various iterations and discussions between all of the authors.
  • Scope and Focus: Studies must address fraud performed online and use AI-based methodologies for analysing online fraudulent activities. The focus was on studies using textual data, whether from scammers, victims, or victim reports, to understand, detect, or analyse online fraud activities. Our goal was to understand the state-of-the-art models designed to prevent and detect scams before the victim gets defrauded. Studies that analysed money transactions, credit card transactions, and cryptocurrency transactions were excluded from this review, as they do not use text data.
  • AI-Methodology: Papers had to include a methodology or similar section where the authors discuss their AI implementation and fine-tuning along with the accuracy of their model. Finally, we considered studies published after 2019.
  • Publication Time Frame: Papers published between January 2019 and March 2024 were included in this review. This period was selected to manage the overwhelmingly large volume of online fraud papers and align with our available resources. Also, we believe that studies conducted before 2019 are less likely to reflect recent advancements in AI methods. Given the rapid evolution of AI-based models, our time frame ensures the inclusion of the most up-to-date and relevant research.